OpenStudy (anonymous):

I have found the Markov Process to be .5 .2 .3 .1 .7 .2 .4 .4 .2 that state 1 is Industry, that state 2 is Small Business, and that state 3 is Self-Employed. How do you find the long-run distribution of people among these various types of employment?

OpenStudy (zarkon):

Find the eigenvector that corresponds to an eigenvalue of 1

OpenStudy (anonymous):

It's also \[\lim_{N \rightarrow \infty}P ^{N}\] where P is that transition matrix. Here is an approximation that might be useful to you in checking your results. 0.28070181, 0.49122801, 0.22807019 0.28070172, 0.49122812, 0.22807017 0.28070177, 0.49122805, 0.22807018 I used the Python numpy library.

OpenStudy (zarkon):

\[\left[\frac{16}{57},\frac{28}{57},\frac{13}{57}\right]\]